OpenStudy (anonymous):

Find the point on the terminal side of θ = pi/4 that has an x coordinate of 2.

OpenStudy (anonymous):

(2,y)

jimthompson5910 (jim_thompson5910):

It may help to draw a picture 

OpenStudy (anonymous):

I don't know how to do this without the hypotenuse

jimthompson5910 (jim_thompson5910):

use the trig identity tan(angle) = opposite/adjacent tan(pi/4) = y/2

OpenStudy (anonymous):

Do I have to make tan(pi/4) into degrees? Sorry for my confusion!

jimthompson5910 (jim_thompson5910):

its ok, make sure you're in radian mode

jimthompson5910 (jim_thompson5910):

do you have a calculator with you?

OpenStudy (anonymous):

http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html this is what I have

jimthompson5910 (jim_thompson5910):

ok, make sure the radio button "Rad" is selected

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

unfortunately, this only supports graphs, but you can still use this as a scientific calculator type in what you see attached

jimthompson5910 (jim_thompson5910):

notice how the y coordinate of each point is 1

jimthompson5910 (jim_thompson5910):

so this means tan(pi/4) = 1

OpenStudy (anonymous):

okay so y would be 2 because 1=y/2 is y=2 right?

jimthompson5910 (jim_thompson5910):

you are 100% correct

OpenStudy (anonymous):

yes! Thank you so much! I thought the line on the calculator meant nothing! :-) Thanks again!

jimthompson5910 (jim_thompson5910):

you're welcome, the calculator is good for graphing, but it may be a bit clunky for other things (like simple arithmetic) That's why I recommend using graph calc (which is what I use) http://www.graphcalc.com/

OpenStudy (anonymous):

Thanks so much! :) and I have one question, how did you know to use tan?

jimthompson5910 (jim_thompson5910):

Because tangent deals with the opposite and adjacent sides The opposite side is 'y' in this case. The adjacent side is 2 in this case.

OpenStudy (anonymous):

Oh that makes total sense! Thanks again!

jimthompson5910 (jim_thompson5910):

Use the mnemonic: SOH CAH TOA SOH: Sine - Opposite over Hypotenuse CAH: Cosine - Adjacent over Hypotenuse TOA: Tangent - Opposite over Adjacent And you can pronounce it in your mind as "soak-a-toe-ah"

OpenStudy (anonymous):

Cool thank you!

jimthompson5910 (jim_thompson5910):

yw